Show your wark to receve fll credt Name Nitrous acid decomposes according to the following reaction How may gams of each of the products can be formed from the decomposition o635fHNO


i need all the answers


i need all the answers


i need all the answers please
Answer #1

1)

3HNO2-> 2NO +HNO3 +H2O

molar mass of HNO2=47
moles of HNO3=63.5/47 =1.351

as per reaction , 3 moles of HNO2 gives 2 moles of NO,1 moles of HNO3 and 1 mole of H20

so,moles of NO=2/3 *1.351=0.90    
grams of NO= 0.90*30=27g

moles of HNO3=1/3*1.351=0.45
grams of HNO3=0.45*63=28.35g

moles of H2O=1/3 *1.351=0.45
grams of H2O=0.45*18=8.1g


2)
The image uploaded by you is not clear ,hence what i make of it is adding now..If there is any change comment and i will give solution.
C3H8 (gas) + 5 O2 (gas) --------> 3 CO2 (gas) + 4 H2O (water) + Heat
2 C7H14 +21 O2 = 14 CO2 + 14 H2O

3)
H2+Cl2 -> 2HCl

moles of HCl=74.30g/36.5 = 2.036
moles of H2 =2.13/2 =1.065

as per equation,1mole H2 gives 2 moles HCl
hence, 1.065 moles give =2.13moles
mass of HCl theoretical =2.13*36.5 = 77.745g

But we only get 2.036

percent yield =actual/calculated *100 = 74.3/77.745 *100 =95.57%